OpenStudy (anonymous):

Find the y-coordinate of the vertex and the axis of symmetry for f(x)=8x-2x^2

OpenStudy (campbell_st):

well thew axis of symmetry for a parabola \[y = ax^2 + bx + c\] is \[x = \frac{-b}{2a}\] in your question b = 8 and a = -2 substitute these into the equation to find the line of symmetry. to find the vertex, substitute the value of x from the line of symmetry into the original equation. The vertex lies on the line of symmetry. hope this helps.